OpenStudy (anonymous):

use implicit differentiation to find equation of tangent line to -1x^2+3xy+4y^3=-8 at the point (-4,2)

OpenStudy (nenadmatematika):

\[-2x+3y+3xy \prime+12y^2y \prime=0\] \[y \prime(3x+12y^2)=2x-3y\] \[y \prime=(2x-3y)/(3x+12y^2)\] if you plug in the given point (-4,2) you get:\[(-8-6)/(-12+48)=-14/36=-7/18\]

OpenStudy (anonymous):

i need it in the form of y=mx + b, i forgot to mention. how do i find b

OpenStudy (bahrom7893):

did i not answer this before?

OpenStudy (anonymous):

bathrom no, i made a mistake with the numbers

OpenStudy (nenadmatematika):

well I think : y-2=-7/18(x+4) so y=-7/18x+4/9 ....I would like someone to check this just in case :D

OpenStudy (anonymous):

*bahrom

OpenStudy (anonymous):

nena thats right, thanks for the help

OpenStudy (nenadmatematika):

haha :D I didn't notice that you wrote ''bathroom'' above haha :D

OpenStudy (anonymous):

typo

OpenStudy (nenadmatematika):

it's OK...you made me laugh so good :D bahrom is nice dude he won't be mad :D

OpenStudy (anonymous):

yeah, he helped me out with this question earlier, but I accidentally gave him a couple different numbers which I just realized so his solution wasn't correct.

OpenStudy (bahrom7893):

no for the numbers you gave me, my solution was correct :P

OpenStudy (anonymous):

yeah. true.

OpenStudy (anonymous):

thats all the help i needed thanks to both